OpenStudy (anonymous):

what's the derivative of 1/(1-7x)2 and can you explain?

OpenStudy (anonymous):

hi welcome to Openstudy! do you know how to apply the quotient rule?

geerky42 (geerky42):

\(\dfrac{1}{(1-7x)^2} = (1-7x)^{-2}\) Apply chain rule from here. Quotient rule is unnecessary here.

OpenStudy (anonymous):

oh yeah, sorry.. i'm thinking about something else

OpenStudy (anonymous):

Thanks so it would be -21/(1-7x)^4 right?

OpenStudy (anonymous):

  simplify...